# Is That Rain? Understanding Effects on Visual Odometry Performance for Autonomous UAVs and Efficient DNN-based Rain Classification at the Edge

### v1

@@ I Introduction
Autonomous Unmanned Aerial Vehicles (UAVs) are set to become central to a variety of industrial applications ranging from first response to infrastructure communications to deliveries, among myriad other UAV-based Internet of Things (IoT) services [1]. In each case, UAVs will require the ability to safely navigate under variable weather conditions. Although much recent research attention has been paid to autonomously navigating complex environments characterised by the presence of obstacles and dynamic disturbances from airflow, little effort has been afforded to understanding the effects of other dynamic environmental factors - particularly rain. Visual odometry (VO) and visual inertial odometry (VIO) leveraging depth cameras are one of the most promising methods to achieve autonomous navigation [2, 3], however, little attention has been paid to understanding the effects of rainy conditions in terms of tracking errors that might be expected when the camera lens is exposed to various rainy conditions.
@@
The most relevant contributions in the literature concerning this or similar problems have emerged from the autonomous driving point of view [4, 5, 6]. In these cases, the images used in the analysis and estimation of the environmental conditions are taken from inside the cockpit; therefore, the water droplets are on the windshield and not directly on the camera lens. On the other hand, such contributions cannot be directly applied to developing reliable autonomous UAV applications because, in such scenarios, the UAVs navigating in rainy conditions may often have the camera lens directly exposed to the environment. As a result, we are motivated to study and analyse the effects of rain on a VO system suitable for autonomous UAVs with direct lens exposure. Given the absence of a suitable relevant dataset, we designed a set of laboratory experiments to simulate a flight at a low altitude (i.e., a scene with objects) under a variety of rainy conditions to collect and label with a view to identifying and classifying precipitation conditions in real time.
@@
  • We describe the training, testing, and comparison of three state-of-the-art DNNs to explore the feasibility of the approach and explore the trade-offs between performance and resources needed  [^2] The source code and all research artefacts will be available open source on acceptance.. Our results demonstrate that various off-the-shelf DNN-architectures developed for mobile or constrained processors offer excellent low-latency classification accuracy under almost all conditions. The best performing model is MobileNetV3 small reaching an accuracy of 90% with a frame rate of 93 FPS and a memory footprint of 1.28 MB.
@@ II Related Works
Considering the variety of under-explored adverse weather conditions, we focus on rain given that it is one of the most common variable conditions, and can be expected to be responsible for significant perturbation to VO systems and consequently damaging to the accuracy and safety of local trajectory planning. Accordingly, we focus on the camera system responsible for the sensor data provided as input to the VO algorithm. Raindrops can bring different perturbations to the sensing and perception system, such as refraction or reflection of light rays. This results in pixel value fluctuation, thus causing the inaccurate processing of the raw images. For instance, DL-based algorithms (e.g., object detection and image classification) are heavily affected by rain, making them unreliable in such conditions. Thus, it is important to be aware of the navigation condition to optimize the involved image processing and computer vision algorithms [28].
@@ III-D Experimental Environment and Settings
The experiments were conducted in a controlled indoor laboratory, which is representative of a challenging and high-entropy scenario. Experiments were designed for two different navigation conditions, namely static and moving. In the static condition, the VO system is stationary in all axes and simulates a drone hovering. In the moving condition, the VO system follows a rectangular trajectory of size $140\times 160$ cm with a fixed altitude. The constant height allows the simulation of the drone’s navigation when it reaches the desired altitude. In this experiment, a specifically trained user moves the VO system over a stool to ensure a constant velocity of around 0.2 m/s (shown in Fig. 1). We use a particularly low velocity to exclude its contribution to the analysis, thus focusing only on the rain effect and perturbation in the navigation system. Moreover, we simulate different rain conditions with a sprayer by changing its distance and inclination from the VO system, taking inspiration from [37]. In particular, we use the rain conditions shown in Table I. In addition, we conducted experiments in clear conditions without simulated rain to have a reference for the other experiments.
Overall, the slanting rain scenario has been designed to show and analyse the effect of raindrops directly on the camera lens. This scenario is the most common, as it is present during navigation at medium/high velocities. In contrast, the vertical rain scenario does not directly show the raindrop effect on the camera lens but happens as a small raindrop aggregation. This scenario reflects hovering or navigation at low velocities. Furthermore, we measured the time required to recover from the slanting rain condition (i.e., the most severe) needed to ensure a VO navigation with an average error below 30 cm, thus ensuring an acceptable error that does not lead to hazardous situations, especially in higher altitude flights.
@@ IV-B Deep Neural Network Specifications
Table II summarizes the number of parameters and the memory footprint of the three architectures. Even though they have a deep structure, they present a low memory footprint because of the innovative blocks that compose them.
@@ V Experimental Results & Analysis
In the first part of this section, we present the experimental results analysed with the setup presented in Section III. We analyse the experiments in static and moving conditions. First, we use the standard deviation to provide the error of the path estimation of the VO system. In the second, we use the root mean square error (RMSE) to provide, on average, the error on the path estimation of the VO system by using the clear condition as the reference. Moreover, we provide the restoring time for the slanting rain scenario needed to ensure navigation with an error below 30 cm. In the second part, we present the results of the DNN test. In particular, we use the confusion matrix of the 7-class classifier to compute average accuracy, precision, recall, and f1-score. These results are used to identify the best performing DNN.
@@ V-A VO System in Static and Moving Conditions
The VO system developed in Section III has been tested in static and moving conditions. For each condition, we evaluated different rain intensities and modalities namely slanting heavy rain, vertical heavy rain, slanting medium rain, vertical medium rain, slanting low rain, and vertical low rain (as shown in Table III).
@@ V-B DNN Test
Moreover, we provide the overall accuracy, precision, recall, and f-score in Table VIII. The computational latency of the three classifiers on the Intel NUC 11 is shown in Table IX. The performance of the three DNNs is comparable, ensuring an accuracy, precision, recall, and f1-score over 90%. On the other hand, MobileNet V3 Small presents the lowest memory footprint (Table II) and the lowest classification latency of around 10 ms (Table IX); thus it is preferable to implement a stable, real-time, and reliable on-the-edge solution in small-size UAVs.
@@
Classifiers Avg. (s) Var.(×$\times 10^{-5}$) ($s^{2}$) Max. (s) Min. (s)
MobileNet V2 0.0706 3.2123 0.0944 0.0651
@@ VI Conclusion
We demonstrate a basis for the development of solutions to mitigate the effects of rain and make the autonomous UAV systems more ‘aware’ of dynamic in situ environmental conditions. Having trained and compared models based on three candidate DNN architectures, we have demonstrated that an accuracy of around 90% can be achieved in classifying the colour images used by the VO system across 7 classes: clear, slanting heavy rain, vertical heavy rain, slanting medium rain, vertical medium rain, slanting low rain, and vertical low rain. The best-performing architecture has reached a frame rate of 97 FPS, achieving approximately real-time processing.

### v2

@@ I Introduction
Autonomous Unmanned Aerial Vehicles (UAVs) are set to become central to a variety of industrial applications ranging from first response to infrastructure communications to deliveries, among myriad other UAV-based Internet of Things (IoT) services [1]. In each case, UAVs will require the ability to safely navigate under variable weather conditions. Although much recent research attention has been paid to autonomously navigating complex environments characterised by the presence of obstacles and dynamic disturbances from airflow [2, 3], little effort has been afforded to understanding the effects of other dynamic environmental factors, particularly rain, as also highlighted in recent articles [4, 5]. Visual odometry (VO) and visual inertial odometry (VIO) leveraging depth cameras are one of the most promising methods to achieve autonomous navigation [6, 7], however, little attention has been paid to understanding the effects of rainy conditions in terms of tracking errors that might be expected when the camera lens is exposed to various rainy conditions.
@@
The most relevant contributions in the literature concerning this or similar problems have emerged from the autonomous driving point of view [8, 9, 10]. In these cases, the images used in the analysis and estimation of the environmental conditions are taken from inside the cockpit; therefore, the water droplets are on the windshield and not directly on the camera lens. On the other hand, such contributions cannot be directly applied to developing reliable autonomous UAV applications because, in such scenarios, the UAVs navigating in rainy conditions may often have the camera lens directly exposed to the environment. While some examples of datasets collected with exposed camera lenses exist, e.g., [11] and [12], these are also in the context of autonomous driving where the instances of ‘rain’ or ‘rainy’ conditions are limited, and the lack of systematic collection and labeling granularity (i.e., within class) renders them insufficient for our purposes. As a result, we are motivated to study and analyse the effects of rain on a VO system suitable for autonomous UAVs with direct lens exposure. Given the absence of a suitable relevant dataset, we designed a set of laboratory experiments to simulate a flight at a low altitude (i.e., a scene with objects) under a variety of rainy conditions to collect and label with a view to identifying and classifying precipitation conditions in real time.
@@
  • We present the training, testing, and comparison of three state-of-the-art DNNs to evaluate the feasibility of the approach and analyze the performance and resource requirements. Our results demonstrate that various off-the-shelf DNN-architectures developed for mobile or constrained processors offer excellent low-latency classification accuracy under almost all conditions. The top-performing model is MobileNetV3 Small achieves 90% accuracy with a frame rate of 93 FPS and a memory footprint of just 1.28 MB.
@@ II Related Works
Considering the variety of under-explored adverse weather conditions, we focus on rain given that it is one of the most common variable conditions, and can be expected to be responsible for significant perturbation to VO systems and consequently damaging to the accuracy and safety of local trajectory planning. Accordingly, we focus on the camera system responsible for the sensor data provided as input to the VO algorithm. Raindrops can bring different perturbations to the sensing and perception system, such as refraction or reflection of light rays. This results in pixel value fluctuation, thus causing the inaccurate processing of the raw images. For instance, DL-based algorithms (e.g., object detection and image classification) are heavily affected by rain, making them unreliable in such conditions. Thus, it is important to be aware of the navigation condition to optimize the involved image processing and computer vision algorithms [34]. Of course, the performance achieved with visual odometry methods can be enhanced via sensor fusion, which may include wireless-inertial methods to cope with extreme weather conditions, low light, or camera saturation. For example, [35] demonstrates how two coupled modules with RF angle-of-arrival (AoA) positioning can estimate UAV attitudes, fused with odometry measurements, to optimize vehicle poses.
@@ III-D Experimental Environment and Settings
The experiments were conducted in a controlled indoor laboratory, which is representative of a challenging and high-entropy scenario. Experiments were designed for two different navigation conditions, namely static and moving. In the static condition, the VO system is stationary in all axes and simulates a drone hovering. In the moving condition, the VO system follows a rectangular trajectory of size $140\times 160$ cm with a fixed altitude. The constant height allows the simulation of the drone’s navigation when it reaches the desired altitude. In this experiment, a specifically trained user moves the VO system over a stool to ensure a constant velocity of around 0.2 m/s (shown in Fig. 1). We use a particularly low velocity to exclude its contribution to the analysis, thus focusing only on the rain effect and perturbation in the navigation system. Moreover, we simulate different rain conditions with a sprayer by changing its distance and inclination from the VO system, taking inspiration from [44]. In particular, we use the rain conditions shown in Table I. In addition, we conducted experiments in clear conditions without simulated rain to have a reference for the other experiments. Overall, the slanting rain scenario has been designed to show and analyse the effect of raindrops directly on the camera lens. This scenario is the most common, as it is present during navigation at medium/high velocities. In contrast, the vertical rain scenario does not directly show the raindrop effect on the camera lens but happens as a small raindrop aggregation. This scenario reflects hovering or navigation at low velocities. Furthermore, we measured the time required to recover from the slanting rain condition (i.e., the most severe) needed to ensure a VO navigation with an average error below 30 cm, thus ensuring an acceptable error that does not lead to hazardous situations, especially in higher altitude flights.
@@ IV-B Deep Neural Network Specifications
Table II summarizes the number of parameters and the memory footprint of the three architectures. Even though they have a deep structure, they present a low memory footprint because of the innovative blocks that compose them. MobileNetV3 Small uses significantly fewer parameters compared to other models, resulting in lower memory usage of only 1.28 MB. This is 24% less than MobileNetV2 and less than half that of SqueezeNet. This reduction in parameters is achieved through a combination of design and architectural optimizations in MobileNetV3 Small.
@@ V Experimental Results & Analysis
In the first part of this section, we present the experimental results analysed with the setup presented in Section III. We analyse the experiments in static and moving conditions. First, we use the standard deviation to provide the error of the path estimation of the VO system. In the second, we use the root mean square error (RMSE) to provide, on average, the error on the path estimation of the VO system by using the clear condition as the reference.
We also report the time required in the rain scenario to ensure a return to navigation accuracy with an acceptable error. We define ‘Restoring Time’ as the duration needed to reduce the error to below 30 cm after encountering adverse rain conditions. In our experiments, the system remained operational during and after the rain emulation. Once the rain ceased, the error remained above 30 cm for some tens of seconds before stabilizing within acceptable limits. The longest Restoring Time, approximately 33 seconds, was observed during slanting rain due to the increased likelihood and intensity of lens wetting. Concerning DNN performance, we analyze the confusion matrix of the 7-class classifier, as provided by the source code, to calculate average accuracy, precision, recall, and F1-score. These metrics are used to identify the best-performing DNN.
@@ V-A VO System in Static and Moving Conditions
The VO system developed in Section III has been tested in static and moving conditions. For each condition, we evaluated different rain intensities and modalities, namely slanting heavy rain, vertical heavy rain, slanting medium rain, vertical medium rain, slanting low rain, and vertical low rain (as shown in Table III).
@@ V-B DNN Test
Moreover, we provide the overall accuracy, precision, recall, and f-score in Table VIII. The computational latency of the three classifiers on the Intel NUC 11 is shown in Table IX. The performance of the three DNNs is comparable, ensuring an accuracy, precision, recall, and f1-score over 90%. On the other hand, MobileNet V3 Small presents the lowest memory footprint (Table II) and the lowest classification latency of around 10 ms (Table IX), that means achieving slightly more than 93 FPS, a marked improvement over other methods.
The update rate or control frequency of an Electronic Speed Controller (ESC) for UAVs is generally in the range of 50 to 500 Hz. This means that the ESC receives and processes control signals from the flight controller approximately 50 to 500 times per second. This update rate allows the ESC to adjust the motor speed rapidly, ensuring precise control of the UAV’s propulsion system [51]. The latency of 10.7ms (93 FPS) achieved by our work, permits taking immediate decisions and intervening promptly on drone control with cycle-accuracy. For this reason, MobileNetV3 Small is a suitable choice for near real-time applications on edge devices such as small-size UAVs.
@@
Classifiers FPS Avg. (ms) Var.(×$\times 10^{-5}$) ($s^{2}$) Max. (s) Min. (s)
MobileNetV2 14.16 70.6 3.2123 0.0944 0.0651
@@ VI Conclusion
We demonstrate a basis for the development of solutions to mitigate the effects of rain and make the autonomous UAV systems more ‘aware’ of dynamic in situ environmental conditions. Having trained and compared models based on three candidate DNN architectures, we have demonstrated that an accuracy of around 90% can be achieved in classifying the colour images used by the VO system across 7 classes: clear, slanting heavy rain, vertical heavy rain, slanting medium rain, vertical medium rain, slanting low rain, and vertical low rain. The best-performing architecture has reached a frame rate of 93 FPS, achieving a latency of $\sim$10 ms, which is sufficient to enable near real-time UAV control.